how do i find the lenght of the three pieces? A 56 inch board is to be divided into 3 pieces so that the second piece is twice as long as the first, and the third piece is 4 times as long as the first. Find the length pf all three pieces.
Found 2 solutions by mananth, Maths68:
Answer by mananth(16946)

first piece =x
second piece = 2x
third piece = 4x
x+2x+4x=56
7x=56
/7
x=8
first piece =8
second piece = 2x=> 16
third piece = 4x=>32
8+16+32=56

Answer by Maths68(1474)

Let
Length of first piece = x
Given
Length of Board = 56 inches
Length of Second piece = 2x
Length of third piece = 4x
Total length of three pieces = Length of Board
x+2x+4x=56
7x=56
x=56/7
x=8

Length of first piece = 8 inches
Length of Second piece = 16 inches
Length of third piece = 32 inches